### v1

# Sparse Model Soups: A Recipe for Improved Pruning via Model Averaging

Abstract
Neural networks can be significantly compressed by pruning, leading to sparse models requiring considerably less storage and floating-point operations while maintaining predictive performance. Model soups (Wortsman et al., 2022a) improve generalization and out-of-distribution performance by averaging the parameters of multiple models into a single one without increased inference time. However, identifying models in the same loss basin to leverage both sparsity and parameter averaging is challenging, as averaging arbitrary sparse models reduces the overall sparsity due to differing sparse connectivities. In this work, we address these challenges by demonstrating that exploring a single retraining phase of Iterative Magnitude Pruning (IMP) with varying hyperparameter configurations, such as batch ordering or weight decay, produces models that are suitable for averaging and share the same sparse connectivity by design. Averaging these models significantly enhances generalization performance compared to their individual components. Building on this idea, we introduce Sparse Model Soups (SMS), a novel method for merging sparse models by initiating each prune-retrain cycle with the averaged model of the previous phase. SMS maintains sparsity, exploits sparse network benefits being modular and fully parallelizable, and substantially improves IMP’s performance. Additionally, we demonstrate that SMS can be adapted to enhance the performance of state-of-the-art pruning during training approaches.
1 Introduction
State-of-the-art Neural Network architectures typically rely on extensive over-parameterization with millions or billions of parameters (Zhang et al., 2016). In consequence, these models have significant memory requirements and the training and inference process is computationally demanding. However, recent work (LeCun et al., 1989; Hassibi and Stork, 1993; Han et al., 2015; Gale et al., 2019; Lin et al., 2020; Hoefler et al., 2021; Zimmer et al., 2022, 2023) has demonstrated that these resource demands can be significantly reduced by pruning the model, i.e., removing redundant structures such as individual parameters or groups thereof. The resulting sparse models demand considerably less storage and floating-point operations (FLOPs) during inference, while retaining performance comparable to dense models.
A different line of research has shown that the performance of a predictor can be significantly enhanced by leveraging multiple models, instead of selecting the best one on a hold-out validation dataset and discarding the rest. Such ensembles combine the predictions of $m\in\mathbb{N}$ individually trained models by averaging their output predictions (Ganaie et al., 2021; Mehrtash et al., 2020; Chandak et al., 2023; Fort et al., 2019). Prediction ensembles have been shown to improve the predictive performance and positively impact predictive uncertainty metrics such as calibration, out-of-distribution generalization as well as model fairness (see e.g. Lakshminarayanan et al., 2017; Mehrtash et al., 2020; Allen-Zhu and Li, 2023; Ko et al., 2023). A significant drawback of ensembling is that all models have to be evaluated during deployment: the inference costs are hence increasing by a factor of $m$, a problem that has partially been addressed by leveraging an ensemble of sparsified, more efficient models (Liu et al., 2021; Whitaker and Whitley, 2022; Kobayashi et al., 2022).
Figure 1: Constructing the average (middle) of two networks with different sparsity pattern (left, right) can reduce the overall sparsity level, turning pruned weights (dashed) into non-zero ones (solid). Weights reactivated in the averaged model are highlighted in orange.
Several works propose to instead average the parameters (Izmailov et al., 2018; Wortsman et al., 2022a; Rame et al., 2022; Matena and Raffel, 2022), constructing a single model for inference. As opposed to prediction ensembles, which require sufficiently diverse models to enhance performance, such Model Soups (Wortsman et al., 2022a) require models that lie in a linearly connected basin of the loss landscape. Training two models from scratch with different random seeds will result in models whose parameter average will perform much worse than the individual models (Neyshabur et al., 2020) with recent research exploring neuron permutation to align them within a single basin (Singh and Jaggi, 2020; Ainsworth et al., 2023). Additionally, the computational advantages of sparse networks cannot be readily utilized, as averaging models with different sparse connectivities reduces the overall sparsity level (cf. Figure 1) and may require to prune again (Yin et al., 2022a, b), potentially resulting in further performance degradation.
In this work, we tackle the challenge of concurrently leveraging sparsity as well as the benefits of combining multiple models into a single one. We draw our inspiration from recent work in the domain of transfer learning (Neyshabur et al., 2020; Wortsman et al., 2022a; Rame et al., 2022), which has shown that fine-tuning multiple copies of a pretrained model, differing only in random seed, yields models sufficiently similar for averaging and sufficiently diverse for performance improvements. At the core of our work lies the observation that a single prune-retrain phase in standard prune after training strategies, such as Iterative Magnitude Pruning (IMP) (Han et al., 2015), closely resembles the transfer learning paradigm. Starting from a pretrained model, the optimization objective shifts abruptly, either due to a new target domain or constraints imposed by the pruning subspace, followed by a subsequent training process termed ‘fine-tuning’, a term frequently used synonymously with ‘retraining’ to recover from pruning (Hoefler et al., 2021).
We find that, akin to the fine-tuning phase in transfer learning, exploring the retraining phase after pruning with various hyperparameter configurations generates models that are amenable to averaging while sharing the same sparse connectivity by design. Such sparse averages exhibit superior performance compared to both their individual counterparts as well as to models retrained $m$ times as long. Taking it a step further, we initiate the subsequent prune-retrain cycle from the averaged model of the previous phase, which remarkably also enhances the performance of the individual retraining runs before averaging again. Our proposed approach, dubbed Sparse Model Soups (SMS), tackles the aforementioned challenges by enabling inference complexity independent of $m$, utilizes pretrained models without requiring training from scratch, preserves the sparsity pattern while leveraging sparse network benefits, and considerably improves IMP’s performance.
Contributions.
To summarize, our contributions can be stated as follows.
  1. 1.
    We demonstrate that pruning a well-trained model and retraining multiple copies with varied hyperparameter like batch ordering, weight decay, or retraining duration and length, produces models suitable for constructing an averaged model which exhibits superior generalization performance compared to its individual components. Importantly, these models maintain the same sparse connectivity of the pruned parent, preserved in their parameter average.
  2. 2.
    We propose Sparse Model Soups (SMS), a novel method for merging sparse models into a single classifier, leveraging the idea of starting each prune-retrain phase of IMP from an averaged model. SMS significantly enhances the performance of IMP in two ways: first, the average improves upon the individual models in terms of generalization performance and, secondly, the models retrained from an average exhibit better performance compared to those retrained from a single model.
  3. 3.
    We extend our findings to the pruning during training domain by incorporating SMS into magnitude-pruning based methods, yielding substantial performance improvements and enhancing their competitiveness in comparison with other state-of-the-art methods that sparsify during training.
Outline.
We start by formally presenting the context and our concept in Section 2. In Section 3, we back up our findings by providing extensive experimental results on image classification, semantic segmentation and neural machine translation architectures and datasets. We review related literature in Section 4 before concluding with a discussion in Section 5.
2 Methodology: Sparse Model Soups
Algorithm 1 Sparse Model Soups
1:Pretrained model $\theta$
2:Sparse model soup
3:for each prune-retrain cycle do
4:     Prune $\theta$
5:     for $i\leftarrow 1$ to $m$ donormal-$\triangleright$ Fully parallelizable
6:         $\theta_{i}\leftarrow\theta$
7:         Retrain $\theta_{i}$ with specific hyperparameters
8:     end for
9:     $\theta\leftarrow$ Merge$(\theta_{1},\ldots,\theta_{m})$
10:end for
11:return $\theta$
Figure 2: Left: Sketch of the algorithm for a single phase and $m=3$. Right: Pseudocode for SMS. Merge()($\cdot$)( ⋅ ) takes $m$ models as input and returns a linear combination of the models (cf. Section 2.2).
In general, we represent neural network parameters as an $n$-dimensional vector $\theta\in\mathbb{R}^{n}$ and use the standard notation $[m]={\{1,\ldots,m\}}$ given $m\in\mathbb{N}$.
2.1 Preliminaries
Our focus lies on model pruning which aims at removing individual weights as exemplified by the previously introduced IMP approach. IMP, a prune after training algorithm, follows a three-stage pipeline. It starts with a pretrained model parameterized by $\theta$, prunes weights with magnitudes below a certain threshold, and then restores predictive power through retraining. This prune-retrain cycle is repeated multiple times, with each pruning step’s threshold determined by the suitable percentile to achieve the desired target sparsity after a predefined number of such phases. Recent studies (Gale et al., 2019; Zimmer et al., 2023) have demonstrated that magnitude pruning results in sparse models with performance competitive to much more complex algorithms.
Given $m$ sparse models $f_{\theta_{i}}$ with weights $\theta_{i}\in\mathbb{R}^{n},i\in[m]$, prediction ensembles construct a model as the functional equivalent of the average of the models’ output (see e.g. Liu et al., 2021; Whitaker and Whitley, 2022; Kobayashi et al., 2022). This ensemble requires $m$ forward passes for evaluation, but maintains the overall sparsity level. In contrast, our focus lies on examining the performance and sparsity of a single model, namely a linear combination of other models. Given scalars $\lambda_{i}\in\mathbb{R},i\in[m]$, we consider the prediction function $f_{\bar{\theta}}$, parameterized by the weights given by
$\bar{\theta}=\sum_{i\in[m]}\lambda_{i}\theta_{i}.$ (1)
A special case occurs when $\lambda_{i}=1/m$ for all $i$, resulting in $\bar{\theta}$ representing the average of all networks. Averaging the weights of arbitrary sparse models can result in reduced overall sparsity, as different networks may possess distinct sparse connectivities, causing the averaging process to eliminate zeros from the tensors (cf. Figure 1). Yin et al. (2022b) and Yin et al. (2022a) address this issue by pruning $\bar{\theta}$ to align with the original networks’ sparsity levels. However, this approach has a notable drawback: if the sparsity patterns differ significantly, pruning-induced performance degradation may occur.
2.2 Sparse Model Soups
Inspired by recent advancements in the transfer learning domain (Neyshabur et al., 2020; Wortsman et al., 2022a), which demonstrate that models fine-tuned from the same pretrained model albeit with different hyperparameter configurations end up in the same loss basin and can be combined into a soup, we hypothesize that a similar behavior can be achieved during fine-tuning or retraining from the same pruned model. Our motivation stems from the resemblance between the transfer learning paradigm and a single phase of IMP. When transitioning from the source to the target domain, the optimization objective changes abruptly, requiring adaptation (i.e., fine-tuning) to minimize the new objective. Similarly, ‘hard’ pruning alters the loss abruptly and requires adapting (i.e., retraining) given the newly added sparsity constraints.
A single phase of this idea is illustrated on the left of Figure 2. The pretrained model $\theta$ is pruned, yielding model $\theta_{p}$, which is then replicated $m$ times. Subsequently, each of the $m$ models is independently retrained with different hyperparameter configurations, such as varying random seeds, weight decay factors, retraining lengths, or learning rate schedules. Finally, the $m$ retrained models are merged into a single model. This process ensures that all $m$ retrained models $\theta_{1},\ldots,\theta_{m}$ share the same sparsity pattern, as they all originate from the same pruned network. However, when combining models after multiple prune-retrain cycles, identical sparsity connectivity between all models is not guaranteed. To address this, we average the models after each phase and commence the subsequent phase with the previously averaged model. The resulting method, termed Sparse Model Soups (SMS), is presented as pseudocode on the right of Figure 2.
SMS offers several benefits and addresses key challenges. First, the inference complexity of the final model remains independent of $m$. The method is highly modular, allowing for different hyperparameter configurations and different $m$ in each phase. Further, the retraining of the $m$ models can be fully parallelized, enhancing efficiency. By initiating each phase with the merged model from the previous one, sparsity patterns are preserved, and the advantages of sparse networks are utilized; as the number of cycles increases, the networks become sparser, potentially leading to further efficiency gains. Moreover, SMS effectively leverages the benefits of large pretrained models without the need for training from scratch.
Merging strategies
In practice, effectively combining models with improved generalization can be challenging. For instance, when using a wide range of weight decay values, models may end up far apart, and merging them could result in decreased performance. Although Equation 1 permits any linear combination of models, we primarily experiment with two convex combination methods proposed by Wortsman et al. (2022a): UniformSoup and GreedySoup. The former assigns an equal factor $\lambda_{i}=1/m$ to each model $i\in[m]$. In contrast, GreedySoup sorts models by their validation accuracy, starting with the best-performing model and progressively adding models from the sorted list only if their inclusion enhances the validation accuracy compared to the previous subset.
3 Experimental results
We first outline our general experimental approach, including datasets, architectures, and metrics. To ensure comparability and standardization in the rapidly evolving field of pruning, we follow the guidelines proposed by Blalock et al. (2020) whenever possible. Our implementation is based on the PyTorch framework (Paszke et al., 2019), using Weights & Biases (Biewald, 2020) for logging. To enable reproducibility, our code is available at github.com/ZIB-IOL/SMS. We evaluate our approach on well-known datasets for image recognition, semantic segmentation, and neural machine translation (NMT), including ImageNet-1K (Russakovsky et al., 2015), CIFAR-10/100 (Krizhevsky et al., 2009), Celeb-A (Liu et al., 2015), CityScapes (Cordts et al., 2016), and WMT16 DE-EN (Bojar et al., 2016). We utilize state-of-the-art architectures, such as ResNets (He et al., 2015), WideResNets (Zagoruyko and Komodakis, 2016), MaxViT (Tu et al., 2022), PSPNet (Zhao et al., 2017), and the T5 transformer (Raffel et al., 2020). We primarily analyze the tradeoff between model sparsity and test performance, using accuracy, mean Intersection over Union (mIoU), and BLEU score as main evaluation metrics. The appendix contains detailed hyperparameters, standard training setups and metric definitions. We also consider the sparsity-induced theoretical speedup (Blalock et al., 2020) and fairness assessed by per-group-recalls as secondary measures. For validation, we use 10% of the training data. Weight magnitude serves as the criterion for unstructured pruning, while for convolutional filters, we follow Li et al. (2016) and prune using filter norms. For retraining, we stick to the linear learning rate schedules proposed by Zimmer et al. (2023).
Recomputing Batch-Normalization (BN) (Ioffe and Szegedy, 2015) statistics is crucial in both pruning and model averaging, as observed by Li et al. (2020) and Jordan et al. (2022), respectively. When reporting test or validation accuracy for single or averaged models, we reset all BN layers and recompute statistics using a forward pass on the entire training dataset. Notably, we fixed the batch order of the dataloader during recomputation, as even random batching could skew results, with more details in the appendix.
3.1 Evaluating Sparse Model Soups
Table 1: WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet: Test accuracy comparison of SMS to several baselines for target sparsities 98% (top) and 90% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-100 (98%)
Sparsity 72.8% (Phase 1) Sparsity 92.6% (Phase 2) Sparsity 98.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 76.50 ±0.16 76.59 ±0.13 76.75 ±0.28 75.55 ±0.60 76.19 ±0.37 76.21 ±0.43 72.67 ±0.29 72.90 ±0.64 73.05 ±0.45
best candidate 75.58 ±0.19 75.71 ±0.08 75.96 ±0.13 74.51 ±0.47 75.01 ±0.74 75.00 ±0.34 71.77 ±0.04 71.77 ±0.37 72.21 ±0.02
mean candidate 75.37 ±0.12 75.58 ±0.03 75.55 ±0.26 74.32 ±0.40 74.71 ±0.48 74.70 ±0.42 71.41 ±0.09 71.61 ±0.40 71.66 ±0.19
SMS (greedy) 76.06 ±0.69 76.43 ±0.24 76.60 ±0.47 75.34 ±0.15 75.39 ±0.44 75.51 ±0.66 72.08 ±0.23 71.86 ±0.64 72.44 ±0.20
best candidate 75.58 ±0.19 75.65 ±0.00 75.94 ±0.15 74.85 ±0.04 74.53 ±0.42 74.57 ±0.21 71.05 ±0.43 71.01 ±0.49 71.47 ±0.23
mean candidate 75.37 ±0.12 75.54 ±0.03 75.54 ±0.27 74.52 ±0.25 74.27 ±0.52 74.20 ±0.31 70.84 ±0.41 70.69 ±0.75 70.87 ±0.01
IMP$+$ 75.85 ±0.26 76.05 ±0.00 75.76 ±0.24 74.09 ±0.24 74.19 ±0.44 74.74 ±0.06 70.92 ±0.07 70.31 ±0.52 71.85 ±0.15
IMP-AVG — N/A — — N/A — 68.19 ±0.44 65.53 ±0.06 63.62 ±0.90
IMP — 75.54 ±0.41 — — 74.09 ±0.13 — — 70.74 ±0.08 —
ImageNet (90%)
Sparsity 53.6% (Phase 1) Sparsity 78.5% (Phase 2) Sparsity 90.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 76.74 ±0.20 76.89 ±0.18 77.01 ±0.05 76.04 ±0.21 76.30 ±0.13 76.49 ±0.12 74.53 ±0.04 74.82 ±0.08 74.96 ±0.16
best candidate 76.07 ±0.01 76.07 ±0.21 76.14 ±0.18 75.48 ±0.16 75.46 ±0.11 75.70 ±0.03 74.00 ±0.03 74.19 ±0.08 74.25 ±0.13
mean candidate 75.99 ±0.04 75.95 ±0.14 75.96 ±0.08 75.40 ±0.11 75.42 ±0.10 75.55 ±0.05 73.94 ±0.03 74.11 ±0.11 74.13 ±0.12
SMS (greedy) 76.74 ±0.19 76.92 ±0.15 76.88 ±0.11 76.12 ±0.18 76.35 ±0.21 76.11 ±0.26 74.58 ±0.03 74.77 ±0.03 74.52 ±0.11
best candidate 76.08 ±0.01 76.08 ±0.21 76.14 ±0.18 75.48 ±0.18 75.53 ±0.24 75.34 ±0.19 74.03 ±0.11 74.21 ±0.00 73.95 ±0.07
mean candidate 75.98 ±0.04 75.95 ±0.14 75.95 ±0.08 75.42 ±0.15 75.45 ±0.21 75.24 ±0.17 73.94 ±0.01 74.09 ±0.03 73.76 ±0.12
IMP+ 76.25 ±0.08 76.21 ±0.14 76.46 ±0.04 75.74 ±0.03 75.87 ±0.11 75.93 ±0.03 74.34 ±0.09 74.56 ±0.24 74.50 ±0.09
IMP-AVG — N/A — — N/A — 72.97 ±0.25 72.58 ±0.01 72.08 ±0.12
IMP — 75.97 ±0.16 — — 75.19 ±0.14 — — 73.59 ±0.04 —
We evaluate SMS against several key baselines, comparing it at each prune-retrain phase to the best-performing single model among all candidates for averaging (best candidate), the mean candidate, regular IMP (i.e., $m=1$) and an extended version of IMP retrained $m$ times as long (IMP+). Unlike SMS, IMP+ is not parallelizable in each phase, with runtime directly dependent on $m$. Moreover, we contrast SMS with executing regular IMP $m$ times, then averaging the models post-final phase (termed IMP-AVG), addressing potential sparsity reduction by re-pruning the model (Yin et al., 2022b). For SMS, we vary random seeds for each of the $m$ models, with additional results in the appendix. In all phases except the first, candidates for averaging, including the best one, are retrained from a previous model soup, distinguishing best candidate from regular IMP, which does not rely on averaging.
Table 1 presents results for three-phase IMP using WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet. The left column lists the baselines, while the three main columns correspond to phases and sparsities, targeting 98% sparsity for CIFAR-100 and 90% for ImageNet. Each main column has three subcolumns, indicating the number of models to average (3, 5, or 10). For full results on different architectures, datasets, target sparsities, and structured pruning, we refer to the appendix. We find that SMS consistently improves upon the test accuracy of the best candidate, often with a 1% or higher margin. This confirms that the models after retraining are averageable, resulting in better generalization than individual models. SMS notably improves upon both classical IMP and its extended retraining variant, IMP+, with up to 2% enhancements even when using $m=3$ splits.
However, remarkably, we find that the best candidates in phases two and three are also consistently better than IMP and often than IMP+, even though they are not part of the average yet. While some improvement is expected when selecting the best model from multiple candidates, the fact that the average test accuracy of all candidates (i.e., mean candidate) also surpasses IMP and IMP+ suggests that starting from a model soup, rather than just a retrained model from the previous phase as classical IMP does, enhances generalization throughout the subsequent phase. This is surprising since the better-performing model soup is more affected by pruning, which further supports the generalization-stability tradeoff (Bartoldson et al., 2020): pruning’s generalization benefit increases with more instability to pruning, as measured by pruning-induced performance degradation. IMP-AVG naively averages models generated by IMP in the final phase. This often leads to reduced sparsity because of differing sparse connectivities (cf. Figure 1), requiring a subsequent pruning that tends to degrade performance compared to individual models.
In summary, we find that averaging after each phase and starting subsequent phases from the soup of the previous phase capitalizes on two dynamics, which often enable significant improvements over IMP: first of all, the model soup consistently improves upon individual soup candidates, demonstrating that pruned and retrained models are indeed averageable, resulting in better generalization performance. Secondly, models retrained from a pruned soup also outperform those following the classical prune-retrain cycle. Pruning a model with higher generalization performance yields better models after retraining, despite experiencing a larger pruning-induced performance drop.
Table 1 also distinguishes between the uniform and greedy selection strategies. When varying only the random seed for each of the $m$ models, none appear to converge to a different basin, making it unnecessary to include only a subset in the soup. We observe that the uniform approach usually outperforms the greedy one in terms of both SMS and the best candidate metric: in the last phase of CIFAR-100, the best and mean uniform candidate substantially outperform the best greedy candidate, indicating that previous greedy soups yield less performant models after retraining.
3.2 Examining Sparse Model Merging
Having demonstrated the advantages of SMS, we now delve deeper into the underlying factors contributing to its success and the cases where it fails.
Exploring parameters beyond random seeds.
In the previous section, we focused on varying the random seed for simplicity. Figure 3 presents a scatter plot for One Shot IMP (70% sparsity) of ResNet-50 trained on ImageNet, comparing the effects of varying the random seed, weight decay strength, retraining duration, and initial learning rate of a linearly decaying schedule. We create eight variations for each setting, selecting equally spaced values around the defaults (details in the appendix). Parameter averages are constructed from two-element pairs of models in the uniform soup setting. The plot displays the test accuracy of the averaged model versus the maximal test accuracy of each pair. In this setting, most averaged models show a net improvement over their individual components, indicating that varying different hyperparameters in the retraining phase of One Shot IMP produces models within the same loss basin. Comparing the different hyperparameters, the most substantial improvement is achieved by varying the random seed, which we attribute to the difficulty of controlling the impact of other parameters: the effect of a different random seed is uncontrollable, while more or less aggressive retraining schedules have more pronounced, specific impacts.
Figure 3: Accuracy of average of two models vs. the maximal individual accuracy. All models are pruned to 70% sparsity (One Shot) and retrained, varying the indicated hyperparameters.
Merging IMP models in later phases.
Throughout our experiments, we observed that in later phases of IMP, assuming regular IMP and not starting from an averaged model as in SMS, sparse connectivities begin to diverge, leading to reduced sparsity when averaging. Specifically, while in the first phase two splits of the models converge to the same loss basin, the next pruning might push them into different subspaces, motivating the combination of individual models into a single one at the end of each phase to ensure that effectively averaging them remains possible. We also experimented with weight alignment, as notably proposed by Ainsworth et al. (2023) and Singh and Jaggi (2020) (see Section 4 for a more detailed account), hoping to permute models at later phases to the same linear subspace. Although we managed to reduce the decrease in sparsity, we were unable to fully recover the original sparsity. This approach can only work if different IMP runs share the same distribution of sparsity among layers (i.e., a specific layer in model 1 must have the same sparsity as the same layer in model 2), which is in general not the case.
Figure 4: WideResNet-20 on CIFAR-100: (a) Accuracy difference between the soup and best performing model after One Shot pruning and retraining. The lines for UniformSoup and GreedySoup show the envelope considering all retraining schedules and durations. (b) Difference between the soup and prolonged IMP version after One Shot pruning to 90% and 98% sparsity. The x-axis shows the retraining length of the $m=3$ models in the soup, while the prolonged IMP version is retrained three times as long. Results are averaged over multiple random seeds with min-max bands indicated.
Exploring limitations: Instability to randomness and recovering it.
Frankle et al. (2020) demonstrate that networks become stable to randomness after a certain amount of pretraining. They specifically analyze the training amount required before splitting up a network into two copies, further trained with different random seeds, such that the final models reside within a linearly connected basin. In that vein, several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) study the stability of IMP with weight rewinding (IMP-WR) in the context of the Lottery Ticket Hypothesis (Frankle and Carbin, 2018). Evci et al. (2022) demonstrate that trained lottery tickets converge to the same basin as IMP-WR solutions, while Paul et al. (2023) establish that successive IMP-WR solutions with varying sparsity levels are linearly mode connected, ensuring that the error remains relatively stable along the linear interpolation between two adjacent solutions. In contrast, our focus lies on retraining networks without rewinding, where the network undergoes pruning-induced perturbations but is not rewound to a previous state. Given the instability analysis of Frankle et al. (2020), it seems plausible that different splits of a pruned network converge to the same basin when the sparsity is low and the learning rate is not drastically increased. Likewise, we anticipate that high pruning levels might reduce stability to randomness.
3(a) compares the test accuracy difference between a soup of $m=5$ models and the best candidate test accuracy. We vary retraining length (5, 20, and 50 epochs) and retraining schedule (LLR and ALLR (Zimmer et al., 2023)), plotting the top-performing configuration for each sparsity. Both UniformSoup and GreedySoup improve over individual models by up to 2%. However, in the high sparsity regime, the soup benefit starts to diminish, even with extended retraining time. Interestingly, UniformSoup collapses, while GreedySoup maintains performance no worse than the best individual model. This indicates that beyond a certain sparsification level, stability to randomness decreases. At least one of the five models diverges from the others, and merging all five leads to reduced performance. Conversely, GreedySoup selects mergeable candidates, possibly relying on just a single model.
3(b) compares a sparse model soup of $m=3$ against a prolonged IMP version (IMP+), retrained three times as long as indicated on the x-axis. We vary the retraining schedule (LLR and ALLR) and the merging method (UniformSoup and GreedySoup), plotting the top-performing configuration. The key observation is that for relatively low sparsity and thus instability (red line), it is more effective to train $m$ models in parallel and average them than to train a single model $m$-times as long, even with short retraining time. For high sparsities like 98% (green line), averaging $m$ models retrained for a short time is not as good as training a single one for longer. However, there is a break-even point at around twelve epochs, where the benefit of retraining a single model longer vanishes, and the $m$ models are trained enough to be mergeable. In iterative pruning, SMS is expected to need fewer retraining epochs per phase to gain model soup benefits due to the gradual increase in sparsity.
Pruning has previously been found to disproportionately affect different subgroups of the data, amplifying unfairness (Hooker et al., 2019, 2020; Paganini, 2020). In the appendix, we further study the capabilities of SMS to mitigate such pruning-induced unfairness. Our findings indicate that SMS has a less severe negative impact on individual subgroups than IMP.
3.3 Improving Pruning during Training algorithms
We extend our findings to magnitude-pruning based methods within the pruning during training domain. These methods, unlike IMP, start with a randomly initialized model and sparsify it during regular training. We focus on GMP (Zhu and Gupta, 2017; Gale et al., 2019), DPF (Lin et al., 2020), and BIMP (Zimmer et al., 2023). GMP employs a pruning schedule to dynamically adjust and update a pruning mask, sparsifying the network throughout training. DPF follows the same schedule but enables error compensation by modifying the update rule to use the pruned model’s gradient while updating the dense parameters. BIMP divides the total training budget into a pretraining phase and multiple IMP cycles for the remaining portion, rendering it closest to IMP. As opposed to others, the three approaches can be easily adapted using SMS. BIMP can incorporate SMS into individual phases as previously discussed. Both GMP and DPF prune at uniformly distributed timesteps throughout training. Consequently, we consider the timeframe between two such steps as a phase, during which we create $m$ copies of the recently pruned model, train them in parallel using different random seeds, and merge them before the next pruning step. Unlike IMP or BIMP, no specific retraining schedule ensures convergence in such a phase; instead, GMP and DPF follow the original learning rate schedule.
Table 2 compares the three methods and their SMS-enhanced versions to other state-of-the-art pruning during training methods, including GSM (Ding et al., 2019), DNW (Wortsman et al., 2019), LC (Carreira-Perpinán and Idelbayev, 2018), and DST (Liu et al., 2020). Ensuring a fair comparison we only consider the dense-to-sparse training paradigm, as opposed to pruning at initialization (Lee et al., 2019; Tanaka et al., 2020) and dynamic sparse training methods (Dettmers and Zettlemoyer, 2019; Evci et al., 2020). We applied LC and GSM to both randomly initialized and pretrained models, selecting the best results for each sparsity, noting the original works only applied these to pretrained models. Further, we experimented with STR (Kusupati et al., 2020), but omitted the results as we were unable to cover the exact sparsity range, being controllable only indirectly through regularization parameters. For a complete overview of the training process and specific hyperparameters, please see the appendix.
We find that incorporating SMS into BIMP, GMP, and DPF consistently improves performance, with BIMP improving the most. We attribute this to the decaying learning rate, which allows for convergence within a phase. Secondly, when compared to other methods, SMS emerges as a simple yet effective way to enhance the competitiveness of magnitude-based pruning methods. However, we note that SMS-enhanced methods increase computational cost by branching into $m$ models between phases, a cost which can be mitigated as individual training can be fully parallelized.
Table 2: ResNet-50 on ImageNet: Comparison of BIMP, GMP and DPF with their SMS-extended variants for goal sparsity levels of 70%, 80%, 90%. Each subcolumn denotes the top-1 accuracy and the theoretical speedup achieved by the method. All results are averaged over multiple seeds and include standard deviations. The best, second best, and third best values are highlighted.
ImageNet
Sparsity 70% Sparsity 80% Sparsity 90%
Method Accuracy Speedup Accuracy Speedup Accuracy Speedup
BIMP+SMS 76.20 ±0.09 2.7 ±0.0 75.76 ±0.11 3.7 ±0.0 74.05 ±0.02 6.1 ±0.0
BIMP 75.62 ±0.02 2.7 ±0.0 75.08 ±0.16 3.7 ±0.0 73.53 ±0.05 6.1 ±0.0
GMP+SMS 75.10 ±0.00 2.7 ±0.0 74.48 ±0.00 3.9 ±0.0 73.12 ±0.02 7.7 ±0.0
GMP 74.55 ±0.07 2.7 ±0.0 73.92 ±0.12 4.0 ±0.0 72.81 ±0.00 7.0 ±0.0
DPF+SMS 76.26 ±0.10 2.7 ±0.0 75.85 ±0.05 3.6 ±0.0 74.31 ±0.00 6.0 ±0.0
DPF 75.74 ±0.02 2.6 ±0.0 75.27 ±0.02 3.6 ±0.0 73.88 ±0.01 5.9 ±0.0
GSM 73.69 ±0.70 2.9 ±0.1 72.75 ±0.62 4.5 ±0.3 70.08 ±0.94 9.5 ±0.8
DNW 75.81 ±0.05 2.5 ±0.0 75.35 ±0.21 3.3 ±0.0 74.24 ±0.12 5.5 ±0.1
LC 75.03 ±0.20 2.4 ±0.0 73.87 ±0.62 3.2 ±0.0 67.57 ±2.71 5.1 ±0.0
DST 72.47 ±0.01 4.1 ±0.0 72.32 ±0.03 9.7 ±0.3 71.35 ±0.09 13.2 ±0.4
4 Related Work
We review the related deep learning literature on prediction ensembling, model averaging and connectivity of solutions, focusing especially on studies that intersect with sparsity-related themes. We refer to Hoefler et al. (2021) and Ganaie et al. (2021) for a comprehensive review of sparsification and prediction ensembling, respectively.
Prediction Ensembling.
Lakshminarayanan et al. (2017) employ Deep Ensembles to enhance predictive uncertainty, improving model calibration and generalization to out-of-distribution data. Huang et al. (2017) use Snapshot Ensembling, averaging model parameters at different snapshots during training and leveraging a cyclic learning rate schedule to explore multiple local minima. Similarly, Fast Geometric Ensembling (Garipov et al., 2018) finds an ensemble within a single training run, while Mehrtash et al. (2020) and Chandak et al. (2023) build ensembles by perturbing parameters and leveraging validation data gradients, respectively. In the context of sparsity, Liu et al. (2021) use dynamic sparse training (Mocanu et al., 2018; Dettmers and Zettlemoyer, 2019; Evci et al., 2020) methods to efficiently generate diverse networks for ensembling. Whitaker and Whitley (2022) create diverse prediction ensembles by pruning randomly and retraining $m$ copies of a pretrained network, while Kobayashi et al. (2022) finetune different subnetworks of a pretrained network.
Model Averaging.
Stochastic Weight Averaging (Izmailov et al., 2018) enhances model generalization by averaging parameters across the SGD optimization trajectory. Wortsman et al. (2022a) and Rame et al. (2022) demonstrate the effectiveness of model soups, achieved by averaging multiple finetuned models with varied hyperparameters, leading to better performance and out-of-domain generalization. The approach most similar to our proposal in Section 3.3 is Late-phase learning (von Oswald et al., 2020), which independently trains and then averages specific parameters, albeit not in combination with pruning. Gueta et al. (2023) focus on fine-tuning language models, revealing a clustering-like behavior in weight space, with regions around closely clustered models containing potentially superior models. Croce et al. (2023) explore model soups of adversarially-robust classifiers, while Choshen et al. (2022) enhance base models by merging multiple finetuned ones. Wortsman et al. (2022b) propose robust fine-tuning through averaging zero-shot and fine-tuned models. The very recent and concurrent work of Jolicoeur-Martineau et al. (2023) introduced PopulAtion Parameter Averaging (PAPA), which, similar to SMS, averages independently trained models regularly to ensure that the models do not become to dissimilar for effective averaging. While their approach does not incorporate the pruning of averaged models, we think that their results can be viewed as further support for the approach of SMS.
Combining model averaging with sparsity, Yin et al. (2022b) suggest an extension to Dynamic Sparse Training (DST) methods. During the last 10% of training, they apply a cyclic learning rate and further pruning to explore different sparsification patterns, yielding multiple ‘cheap tickets’. These networks are then integrated into a single model, which, despite densifying upon averaging, is pruned to maintain the initial sparsity level. Yin et al. (2022a) proposes averaging adjacent subnetworks of IMP to create superior lottery tickets, with sparsity reduction counteracted by magnitude pruning. Stripelis et al. (2022) introduce a Federated Learning algorithm, FedSparsify, where a global mask is maintained centrally and updated with local client prunings, with disparities resolved by majority voting. Concurrently and independently from our work, Jaiswal et al. (2023) obtain subsequent pruning masks early by exploring and averaging several mask configurations.
Mode Connectivity.
Neyshabur et al. (2020) demonstrate that training two models from scratch leads to models not linearly connected, while models finetuned from a pretrained model tend to be similar and reside within the same loss basin. Entezari et al. (2022) conjecture that two models trained with different seeds are linear mode connected up to permutations of the network’s neurons and filters. Demonstrating this linear mode connectivity, Ainsworth et al. (2023) propose algorithms to find permutations transforming models into a shared loss basin. Singh and Jaggi (2020) leveraged model fusion for neuron soft-alignment across models and improved convolutional filter pruning by fusing the dense model into the sparse one. In that vein, Benzing et al. (2022) propose a permutation algorithm and surprisingly demonstrate that models also lie in the same loss valley (up to permutation) at time of their initialization. Finally, Jordan et al. (2022) explore ‘variance collapse’ in pre-activations when interpolating deep networks, proposing mitigation strategies. There have been several works (Frankle et al., 2020; Evci et al., 2022; Paul et al., 2023) studying IMP’s stability to randomness, we refer to Section 3.2 for a detailed account.
5 Discussion
Efficient, high-performing sparse networks are crucial in resource-constrained environments. However, sparse models cannot easily leverage the benefits of parameter averaging. We addressed this issue proposing SMS, a technique that merges models while preserving sparsity, substantially enhancing IMP and outperforming multiple baselines. By integrating SMS into magnitude-pruning methods during training, we elevated their performance and competitiveness. Despite the focus on pruning, a single type of network compression, we think that our work serves as an important step towards understanding and improving sparsification algorithms. A potential limitation of SMS may be the computational cost, demanding multiple, yet parallelizable, training runs, which might be restrictive in some settings.
Appendix A Technical details and training settings
Training settings and metrics used.
Table 3 shows the exact pretraining settings for each dataset-architecture pair, reporting the number of epochs used for pretraining, the batch size, weight decay as well as the learning rate used. We kept momentum at the default value of 0.9. The last column reports the performance we achieve when performing regular dense training. We stick to SGD as the optimizer, noting that a variety of other optimization methods for training deep neural networks exist [see e.g. Kingma and Ba, 2014, Pokutta et al., 2020]. For image classification tasks, we report the top-1 test accuracy being the fraction of correctly classified test samples. For semantic segmentation, we used pretrained backbones and evaluated the mean Intersection-over-Union (IoU) on the validation dataset. For NMT, we report the BLEU score on the test set with sequence length limited to 128. We utilized label smoothing and gradient clipping for MaxViT. If needed, we report the theoretical speedup, a metric indicating the FLOPs ratio for inference between dense and sparse models [Blalock et al., 2020]. The speedup, defined as $F_{d}/F_{s}$ where $F_{d}$ and $F_{s}$ are the FLOPs required for dense and pruned models respectively, depends solely on the distribution of pruned weights, not on the values attained by non-zero parameters. FLOPs are computed using a single test set batch, with code adapted from the ShrinkBench framework [Blalock et al., 2020].
Table 3: Exact training configurations in our experiments.
Dataset Network (number of weights) Epochs Batch size Weight decay Learning rate ($t$ = training epoch) Unpruned test accuracy/IoU/BLEU
CIFAR-10 ResNet-18 (11 Mio) 200 128 5e-4 $\eta_{t}=\begin{cases}0.1&t\in[1,90],\\ 0.01&t\in[91,180],\\ 0.001&t\in[181,200]\end{cases}$ 95.0% ±0.04%
CIFAR-100 WRN-20 (26 Mio) 200 128 2e-4 linear from 0.1 to 0 76.5% ±0.1%
ImageNet ResNet-50 (26 Mio) 90 256 1e-4 linear from 0.1 to 0 76.12% ±0.01%
ImageNet MaxViT (31 Mio) 200 256 1e-5 $\eta_{t}=\begin{cases}0.2\frac{t}{20}&t\in[1,20],\\ 0.2&t\in[20,60],\\ 0.02&t\in[61,120],\\ 0.002&t\in[121,160],\\ 0.0002&t\in[161,200]\\ \end{cases}$ 78.0% ±0.02%
CityScapes PSPNet (68 Mio) 300 12 1e-5 $\eta_{t}=\begin{cases}0.1\frac{t}{20}&t\in[1,20],\\ 0.1&t\in[20,100],\\ 0.01&t\in[101,200],\\ 0.001&t\in[201,270],\\ 0.0001&t\in[271,290]\\ 0.00001&t\in[291,300]\\ \end{cases}$ 58.3 IoU ±0.5
WMT16 (EN-DE) T5-small (77 Mio) 5 16 1e-5 $\eta_{t}=\begin{cases}0.1t&t\in[0,1.0],\\ 0.1&t\in[1,2[,\\ 0.01&t\in[2,3[,\\ 0.001&t\in[3,4[,\\ 0.0001&t\in[4,5]\\ \end{cases}$ 24.56 BLEU ±0.007
Pruning settings.
Identifying which weights to remove is essential for successful magnitude pruning, with multiple methods developed to address this. Zhu and Gupta [2017] presented the Uniform allocation that prunes each layer to the same relative sparsity level. Gale et al. [2019] refined this approach to Uniform+, leaving the first convolutional layer dense and capping pruning in the final fully-connected layer at 80%. Evci et al. [2020] reformulate the Erdős-Rényi kernel (ERK) [Mocanu et al., 2018] to consider layer and kernel dimensions for layerwise sparsity distribution. Lee et al. [2020] suggested Layer-Adaptive Magnitude-based Pruning (LAMP), which minimizes output distortion at pruning time measured by the $L_{2}$-distortion on the worst-case input. Throughout this work, we stick to the Global allocation, in which all trainable parameters treated as a single vector and a global threshold is computed to remove parameters independent of the layer they belong to. In experiments where we prune convolutional filters instead of weights, we adopt the $L_{2}$-norm criterion from Li et al. [2016], ensuring a uniform sparsity distribution across layers.
We follow the recommendations of Evci et al. [2020] and Dettmers and Zettlemoyer [2019] and refrain from pruning biases and batch-normalization parameters, as their negligible weight contribution is offset by their significant performance impact. Moreover, for GMP experiments, we opt for the global selection criterion due to its superior performance compared to Uniform+ [Zimmer et al., 2023].
Model Soups and Batch-Normalization statistics.
Throughout our experiments, we also explored variants of LearnedSoup [Wortsman et al., 2022a], which learns the coefficients $\lambda_{i}$ to maximize the validation accuracy. Specifically, we observed improvements when utilizing knowledge distillation techniques [Hinton et al., 2015] with the original pretrained model as the teacher, instead of minimizing the validation loss as suggested by Wortsman et al. [2022b]. Nevertheless, these improvements were marginal, so we opted for UniformSoup and GreedySoup for simplicity and to avoid introducing new hyperparameters.
When constructing an averaged model or changing the parameters in any other way, the Batch-Normalization statistics have to be updated, which can be done by performing a forward pass on (part of) the training data without backpropagation. Since doing so only for model soups could potentially skew the results, we decided to recompute the statistics for single models as well to have a comparable setting independent of whether we changed the parameters or not. In particular, we enforce using the entire train data loader and we fixed its random batch ordering to ensure reproducibility and to avoid the batch ordering having any influence.
Hyperparameters for Pruning during Training algorithms.
Unless stated otherwise, we set weight decay to 1e-4 and momentum to 0.9, with all methods following a linear learning rate schedule starting from 1e-1. For GSM and LC, we select the best result either from scratch or using a pretrained model, applying these methods for 10, 20, or 40 epochs. When extending BIMP, GMP and DPF with SMS, we choose the number of copies to create within a phase, $m$, between 2 and 3. Further, we tuned the epoch at which we begin to train multiple copies between 50 and 75. Otherwise, we applied the following hyperparameter grids.
  • BIMP
    • Initial training budget epochs: $60,75$.
    • Number of pruning phases of equal length: $1,2,3$.
  • GMP
    • Equally distributed pruning steps: $5,9,18,45$.
  • DPF
    • Equally distributed pruning steps: $9,18$.
  • GSM
    • Momentum: $0.9,0.95$.
    • Weight decay: 1e-4, 1e-5.
  • LC
    • Weight decay: 1e-4, 1e-5.
  • DNW
    • Weight decay: 1e-4, 1e-5.
  • DST
    • Weight decay: 1e-4, 1e-5.
    • $\alpha$: 1e-7, 5e-7, 1e-6, 2e-6, 5e-6, 8e-6, 1e-5, 1e-4.
Appendix B Full results
This section contains additional tables and plots. The subsections follow the same structure as the main experimental section.
B.1 Evaluating Sparse Model Soups
Table 4: WideResNet-20 on CIFAR-100 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 90% (top) and 98% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-100 (90%)
Sparsity 53.6% (Phase 1) Sparsity 78.5% (Phase 2) Sparsity 90.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 76.30 ±0.07 76.38 ±0.29 76.45 ±0.34 76.55 ±0.17 76.89 ±0.42 76.94 ±0.25 75.98 ±0.43 76.26 ±0.76 76.67 ±0.02
best candidate 76.03 ±0.06 76.18 ±0.03 76.13 ±0.35 75.88 ±0.49 75.78 ±0.35 75.82 ±0.21 75.15 ±0.05 75.23 ±0.38 75.52 ±0.30
mean candidate 75.88 ±0.00 75.86 ±0.15 75.93 ±0.27 75.50 ±0.21 75.48 ±0.12 75.45 ±0.13 74.99 ±0.22 75.06 ±0.42 75.15 ±0.16
SMS (greedy) 76.28 ±0.10 76.04 ±0.16 76.12 ±0.32 76.16 ±0.21 76.45 ±0.61 76.45 ±0.28 75.48 ±0.16 75.91 ±0.08 75.81 ±0.13
best candidate 76.05 ±0.03 76.18 ±0.03 76.11 ±0.33 75.44 ±0.11 75.66 ±0.19 75.59 ±0.35 75.14 ±0.06 75.08 ±0.35 74.97 ±0.11
mean candidate 75.93 ±0.02 75.86 ±0.15 75.92 ±0.26 75.26 ±0.11 75.34 ±0.15 75.24 ±0.19 74.87 ±0.01 74.81 ±0.23 74.72 ±0.21
IMP$+$ 76.30 ±0.42 75.97 ±0.30 76.27 ±0.02 75.69 ±0.43 75.86 ±0.48 75.91 ±0.25 74.59 ±0.61 74.73 ±0.42 75.00 ±0.57
IMP-AVG — N/A — — N/A — 75.70 ±0.59 75.68 ±0.25 75.60 ±0.23
IMP — 75.64 ±0.21 — — 75.51 ±0.52 — — 74.91 ±0.71 —
CIFAR-100 (98%)
Sparsity 72.8% (Phase 1) Sparsity 92.6% (Phase 2) Sparsity 98.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 76.50 ±0.16 76.59 ±0.13 76.75 ±0.28 75.55 ±0.60 76.19 ±0.37 76.21 ±0.43 72.67 ±0.29 72.90 ±0.64 73.05 ±0.45
best candidate 75.58 ±0.19 75.71 ±0.08 75.96 ±0.13 74.51 ±0.47 75.01 ±0.74 75.00 ±0.34 71.77 ±0.04 71.77 ±0.37 72.21 ±0.02
mean candidate 75.37 ±0.12 75.58 ±0.03 75.55 ±0.26 74.32 ±0.40 74.71 ±0.48 74.70 ±0.42 71.41 ±0.09 71.61 ±0.40 71.66 ±0.19
SMS (greedy) 76.06 ±0.69 76.43 ±0.24 76.60 ±0.47 75.34 ±0.15 75.39 ±0.44 75.51 ±0.66 72.08 ±0.23 71.86 ±0.64 72.44 ±0.20
best candidate 75.58 ±0.19 75.65 ±0.00 75.94 ±0.15 74.85 ±0.04 74.53 ±0.42 74.57 ±0.21 71.05 ±0.43 71.01 ±0.49 71.47 ±0.23
mean candidate 75.37 ±0.12 75.54 ±0.03 75.54 ±0.27 74.52 ±0.25 74.27 ±0.52 74.20 ±0.31 70.84 ±0.41 70.69 ±0.75 70.87 ±0.01
IMP$+$ 75.85 ±0.26 76.05 ±0.00 75.76 ±0.24 74.09 ±0.24 74.19 ±0.44 74.74 ±0.06 70.92 ±0.07 70.31 ±0.52 71.85 ±0.15
IMP-AVG — N/A — — N/A — 68.19 ±0.44 65.53 ±0.06 63.62 ±0.90
IMP — 75.54 ±0.41 — — 74.09 ±0.13 — — 70.74 ±0.08 —
Table 5: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsity 98% given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-10 (98%)
Sparsity 72.8% (Phase 1) Sparsity 92.6% (Phase 2) Sparsity 98.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 95.41 ±0.00 95.46 ±0.13 95.64 ±0.01 95.45 ±0.01 95.59 ±0.09 95.61 ±0.03 94.91 ±0.16 95.24 ±0.11 95.40 ±0.10
best candidate 94.84 ±0.10 94.96 ±0.13 94.94 ±0.07 94.92 ±0.07 95.11 ±0.06 95.12 ±0.02 94.40 ±0.02 94.69 ±0.16 94.81 ±0.09
mean candidate 94.70 ±0.13 94.82 ±0.00 94.75 ±0.08 94.85 ±0.04 94.93 ±0.03 94.94 ±0.01 94.32 ±0.12 94.57 ±0.16 94.66 ±0.13
SMS (greedy) 95.42 ±0.04 95.45 ±0.18 95.50 ±0.05 95.46 ±0.08 95.27 ±0.24 95.32 ±0.20 95.01 ±0.08 94.92 ±0.04 94.94 ±0.25
best candidate 94.84 ±0.10 94.96 ±0.13 94.94 ±0.07 94.96 ±0.26 94.86 ±0.09 94.92 ±0.08 94.55 ±0.11 94.48 ±0.01 94.54 ±0.06
mean candidate 94.70 ±0.13 94.82 ±0.00 94.75 ±0.08 94.80 ±0.15 94.76 ±0.11 94.79 ±0.07 94.44 ±0.06 94.34 ±0.01 94.29 ±0.06
IMP+ 95.18 ±0.08 95.16 ±0.16 95.19 ±0.18 95.02 ±0.11 95.11 ±0.18 95.20 ±0.02 94.62 ±0.28 94.61 ±0.02 94.59 ±0.23
IMP-AVG — N/A — — N/A — 94.44 ±0.28 94.24 ±0.13 93.62 ±0.16
IMP — 94.71 ±0.08 — — 94.92 ±0.01 — — 94.17 ±0.04 —
Table 6: ResNet-18 on CIFAR-10 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 80%, 90%, 95% in the One Shot setting and a retrain length of 20 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-10
Sparsity 80.0% (One Shot) Sparsity 90.0% (One Shot) Sparsity 95.0% (One Shot)
Accuracy of $m=3$ $m=3$ $m=3$
SMS (uniform) 95.49 ±0.05 95.42 ±0.06 95.03 ±0.15
best candidate 95.04 ±0.06 94.84 ±0.03 94.63 ±0.06
mean candidate 94.92 ±0.08 94.76 ±0.03 94.48 ±0.05
SMS (greedy) 95.36 ±0.16 95.41 ±0.07 95.04 ±0.19
best candidate 95.04 ±0.06 94.84 ±0.03 94.63 ±0.06
mean candidate 94.92 ±0.08 94.76 ±0.03 94.48 ±0.05
IMP+ 95.17 ±0.17 95.02 ±0.01 94.72 ±0.24
IMP 95.02 ±0.05 94.71 ±0.28 94.38 ±0.02
Table 7: MaxViT on ImageNet (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 75%, 80%, 85% in the One Shot setting and a retrain length of 10 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
ImageNet
Sparsity 75.0% (One Shot) Sparsity 80.0% (One Shot) Sparsity 85.0% (One Shot)
Accuracy of $m=3$ $m=3$ $m=3$
SMS (uniform) 78.31 ±0.21 78.12 ±0.17 77.59 ±0.24
best candidate 78.11 ±0.07 77.85 ±0.10 77.37 ±0.02
mean candidate 77.83 ±0.01 77.67 ±0.07 77.17 ±0.04
SMS (greedy) 78.21 ±0.14 78.06 ±0.01 77.45 ±0.13
best candidate 78.11 ±0.07 77.85 ±0.10 77.37 ±0.02
mean candidate 77.83 ±0.01 77.66 ±0.06 77.17 ±0.04
IMP$+$ 78.17 ±0.15 77.99 ±0.12 77.68 ±0.23
IMP 78.07 ±0.09 77.88 ±0.06 77.34 ±0.12
Table 8: PSPNet on Cityscapes (unstructured pruning): Validation accuracy comparison of SMS to several baselines for target sparsity 90% given two prune-retrain cycles. We only report results for UniformSoup since we evaluate on the validation dataset. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CityScapes (90%)
Sparsity 68.3% (Phase 1) Sparsity 90.0% (Phase 2)
Accuracy of $m=3$ $m=5$ $m=3$ $m=5$
SMS (uniform) 58.61 ±0.22 58.96 ±0.35 59.28 ±0.73 59.01 ±0.35
best candidate 58.65 ±0.69 58.91 ±0.56 58.83 ±0.54 58.65 ±0.34
mean candidate 58.33 ±0.28 58.35 ±0.30 58.47 ±0.23 58.28 ±0.36
IMP+ 58.76 ±0.18 58.37 ±0.53 58.52 ±0.39 58.63 ±0.38
IMP-AVG — N/A — 58.79 ±0.40 59.01 ±0.07
IMP — 58.40 ±0.28 — — 58.41 ±0.09 —
Table 9: PSPNet on Cityscapes (unstructured pruning): Validation accuracy comparison of SMS to several baselines for target sparsities 60%, 70%, 80% and 90% in the One Shot setting and a retrain length of 10 epochs. We only report results for UniformSoup since we evaluate on the validation dataset. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CityScapes
Sparsity 60.0% (One Shot) Sparsity 70.0% (One Shot) Sparsity 80.0% (One Shot) Sparsity 90.0% (One Shot)
Accuracy of $m=3$ $m=3$ $m=3$ $m=3$
SMS (uniform) 58.85 ±0.46 58.04 ±0.84 57.80 ±0.20 57.61 ±0.51
best candidate 58.61 ±0.86 58.00 ±0.84 57.37 ±0.14 57.45 ±0.44
mean candidate 58.17 ±0.69 57.69 ±0.72 57.21 ±0.16 56.91 ±0.35
IMP$+$ 58.08 ±0.03 58.41 ±0.22 57.69 ±0.23 57.36 ±0.06
IMP 58.09 ±0.57 57.59 ±0.62 57.42 ±0.44 56.94 ±0.54
Table 10: T5 on WMT16 (unstructured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 50%, 60%, 70% in the One Shot setting. We report results using UniformSoup merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
WMT-16
Sparsity 50.0% (One Shot) Sparsity 60.0% (One Shot) Sparsity 70.0% (One Shot)
Accuracy of $m=3$ $m=3$ $m=3$
SMS (uniform) 25.47 ±0.52 25.09 ±0.00 24.51 ±0.43
best candidate 25.39 ±0.03 24.96 ±0.26 24.12 ±0.01
mean candidate 25.16 ±0.08 24.79 ±0.19 24.03 ±0.01
IMP+ 25.36 ±0.12 25.09 ±0.05 24.00 ±0.04
IMP 25.15 ±0.20 24.90 ±0.20 24.04 ±0.28
Table 11: WideResNet-20 on CIFAR-100 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 60% (top) and 80% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-100 (60%)
Sparsity 26.2% (Phase 1) Sparsity 45.6% (Phase 2) Sparsity 60.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 76.54 ±0.19 76.56 ±0.21 76.82 ±0.19 75.99 ±0.07 76.23 ±0.35 76.46 ±0.16 75.34 ±0.11 75.44 ±0.11 75.74 ±0.18
best candidate 75.12 ±0.04 75.19 ±0.06 75.23 ±0.07 74.73 ±0.15 74.90 ±0.12 74.85 ±0.16 74.31 ±0.42 74.45 ±0.11 74.53 ±0.16
mean candidate 74.98 ±0.12 74.87 ±0.06 74.86 ±0.05 74.49 ±0.14 74.53 ±0.14 74.53 ±0.06 73.84 ±0.36 73.95 ±0.19 73.95 ±0.05
SMS (greedy) 76.55 ±0.41 76.45 ±0.24 76.40 ±0.18 75.85 ±0.06 76.10 ±0.38 76.03 ±0.43 75.10 ±0.12 75.24 ±0.20 74.48 ±0.42
best candidate 75.09 ±0.01 75.19 ±0.06 75.21 ±0.10 74.88 ±0.16 74.92 ±0.11 74.80 ±0.23 74.04 ±0.15 74.28 ±0.37 74.23 ±0.22
mean candidate 74.96 ±0.09 74.87 ±0.06 74.88 ±0.02 74.62 ±0.12 74.56 ±0.28 74.43 ±0.20 73.85 ±0.05 73.78 ±0.12 73.87 ±0.14
IMP+ 75.55 ±0.04 75.72 ±0.18 75.92 ±0.16 74.84 ±0.02 75.12 ±0.09 75.68 ±0.47 74.28 ±0.05 74.63 ±0.32 74.73 ±0.94
IMP-AVG — N/A — — N/A — 74.65 ±0.63 75.54 ±0.28 75.49 ±0.33
IMP — 74.96 ±0.20 — — 74.09 ±0.05 — — 73.47 ±0.04 —
CIFAR-100 (80%)
Sparsity 41.5% (Phase 1) Sparsity 65.8% (Phase 2) Sparsity 80.0% (Phase 3)
Accuracy of $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$ $m=3$ $m=5$ $m=10$
SMS (uniform) 75.94 ±0.01 75.99 ±0.40 76.19 ±0.40 74.18 ±0.02 74.23 ±0.27 74.76 ±0.06 71.56 ±0.16 71.59 ±0.14 71.78 ±0.25
best candidate 74.65 ±0.29 74.65 ±0.11 74.78 ±0.17 73.27 ±0.26 73.22 ±0.27 73.71 ±0.18 70.61 ±0.11 70.58 ±0.50 70.96 ±0.33
mean candidate 74.44 ±0.17 74.37 ±0.16 74.39 ±0.13 72.90 ±0.08 72.94 ±0.23 73.19 ±0.09 70.50 ±0.08 70.31 ±0.52 70.40 ±0.23
SMS (greedy) 75.87 ±0.10 75.97 ±0.44 76.14 ±0.52 74.13 ±0.23 74.21 ±0.10 74.48 ±0.38 71.63 ±0.25 71.70 ±0.27 71.06 ±0.87
best candidate 74.70 ±0.23 74.65 ±0.11 74.80 ±0.15 73.18 ±0.22 73.40 ±0.32 73.59 ±0.25 70.86 ±0.30 70.93 ±0.04 70.37 ±0.79
mean candidate 74.47 ±0.13 74.37 ±0.16 74.38 ±0.14 72.95 ±0.33 73.03 ±0.07 73.13 ±0.10 70.56 ±0.45 70.37 ±0.04 69.83 ±0.88
IMP+ 75.06 ±0.04 75.48 ±0.16 75.39 ±0.09 73.44 ±0.22 73.70 ±0.06 73.96 ±0.54 72.06 ±0.10 72.15 ±0.58 72.32 ±0.58
IMP-AVG — N/A — — N/A — 70.52 ±0.18 68.60 ±2.94 69.89 ±1.34
IMP — 73.95 ±0.08 — — 72.71 ±0.15 — — 69.88 ±0.50 —
Table 12: ResNet-18 on CIFAR-10 (structured pruning): Test accuracy comparison of SMS to several baselines for target sparsities 40%, 50%, 60% in the One Shot setting and a retrain length of 20 epochs. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
CIFAR-10
Sparsity 40.0% (One Shot) Sparsity 50.0% (One Shot) Sparsity 60.0% (One Shot)
Accuracy of $m=3$ $m=3$ $m=3$
SMS (uniform) 94.91 ±0.00 94.68 ±0.02 94.32 ±0.10
best candidate 94.42 ±0.04 94.28 ±0.01 93.81 ±0.06
mean candidate 94.31 ±0.01 94.16 ±0.03 93.77 ±0.06
SMS (greedy) 94.87 ±0.02 94.70 ±0.08 94.31 ±0.11
best candidate 94.42 ±0.04 94.28 ±0.01 93.81 ±0.06
mean candidate 94.31 ±0.01 94.17 ±0.04 93.77 ±0.06
IMP+ 94.77 ±0.01 94.38 ±0.23 94.34 ±0.16
IMP 94.45 ±0.18 94.06 ±0.17 93.96 ±0.34
B.2 Examining Sparse Model Merging
Exploring parameters beyond random seeds.
Similar to Figure 3, Figure 5 visualizes the effects of different hyperparameters given One Shot IMP (90% sparsity) with WRN-20 trained on CIFAR-100. Again, we created eight variations for each parameter (random seed, weight decay strength, retraining duration, and initial learning rate in a linear decay schedule) based on equidistant values around the defaults. The scatter plot compares the test accuracy of models in the uniform soup setting (averaged from pairs of models) against the maximum test accuracy within each pair.
Figure 5: WideResNet-20 on CIFAR-100: Accuracy of average of two models vs. the maximal individual accuracy. All models are pruned to 90% sparsity (One Shot) and retrained, varying the indicated hyperparameters.
Exploring limitations: Instability to randomness and recovering it.
Figure 6: ResNet-18 on CIFAR-10: (a) Accuracy difference between the soup and best performing model after One Shot pruning and retraining. The lines for UniformSoup and GreedySoup show the envelope considering all retraining schedules and durations. (b) Difference between the soup and prolonged IMP version after One Shot pruning to 90% and 98% sparsity. The x-axis shows the retraining length of the $m=3$ models in the soup, while the prolonged IMP version is retrained three times as long. Results are averaged over multiple random seeds with min-max bands indicated.
Reducing compression-induced unfairness.
Classification model performance, usually measured by the top-1 accuracy, can mask the disproportionate effect of compression on individual class performance [Hooker et al., 2019, 2020, Paganini, 2020]. Pruning often sacrifices difficult samples, benefiting well-performing classes and worsening the performance of others [Tran et al., 2022]. Recent research by Ko et al. [2023] highlights the benefits of prediction ensembling in enhancing fairness metrics, including minority group performance.
We investigate whether model averaging mitigates pruning’s adverse effects on fairness using ResNet-18 trained on the Celeb-A facial attribute recognition dataset, a fairness benchmark due to its strong sub-group and target label correlation. Table 13 compares the dense model, One Shot SMS, and IMP at sparsities of 90%, 95%, and 97%, where 20 epochs of retraining yields similar top-1 test accuracies for SMS and IMP. We report the recall for the disjoint sub-groups: Male-Blonde (MB), Male-Non-Blonde (MN), Female-Blonde (FB), and Female-Non-Blonde (FN), in addition to top-1 test accuracy. Although SMS and IMP reach similar test accuracy, differences in subgroup performance arise for sparsities above 90%. SMS notably increases recall for MB and FB, the most challenging classes in the dense model. This emphasizes the negative effects of regular pruning through IMP, which sacrifices weakly represented subgroups to maintain high overall accuracy. In contrast, SMS has a less significant impact on MB and FB but leads to a more nuanced decline in easier-to-classify subgroups.
Table 13: ResNet-18 on Celeb-A: Comparison of the pretrained (i.e. dense) base model against SMS and IMP for different sparsity levels. We indicate the top-1 test accuracy as well as the recall on the four different sub-groups. All results are averaged over multiple random seeds with standard deviation included.
Celeb-A
Sub-group Recall
Setting Sparsity Top-1 acc. Balanced acc. MB MN FB FN
Pretrained 0% 98.99 ±0.01 98.09 ±0.02 94.41 ±0.20 99.93 ±0.00 98.48 ±0.00 99.54 ±0.00
SMS 90% 99.02 ±0.00 98.13 ±0.04 94.52 ±0.20 99.92 ±0.00 98.57 ±0.02 99.50 ±0.04
IMP 90% 98.99 ±0.02 98.12 ±0.03 94.52 ±0.20 99.93 ±0.01 98.51 ±0.04 99.52 ±0.03
SMS 95% 98.91 ±0.01 98.08 ±0.05 94.74 ±0.20 99.88 ±0.01 98.53 ±0.07 99.17 ±0.06
IMP 95% 98.79 ±0.01 97.89 ±0.08 94.02 ±0.31 99.92 ±0.01 98.20 ±0.01 99.41 ±0.01
SMS 97% 97.79 ±0.08 96.04 ±0.53 89.44 ±1.78 99.67 ±0.07 97.38 ±0.24 97.68 ±0.49
IMP 97% 97.74 ±0.08 95.73 ±0.32 87.42 ±1.17 99.88 ±0.01 96.71 ±0.07 98.90 ±0.04